OpenStudy (anonymous):

verify the functions are inverses: 1. f(x)=x+7 g(x)=x-7 2. f(x)=2x+1 g(x)=1/2x-1/2 3. f(n)= (-16+n)/4 g(n)=4n+16 4. f(x)=x^2-9 g(x)=x+3

OpenStudy (anonymous):

compose them and see that you get \(x\) back for example \[f\cir g(x)=f(g(x))=f(x-7)=x-7+7=x\] for the first one

OpenStudy (anonymous):

@satellite73 I got x for both

OpenStudy (anonymous):

can anyone help with #3 thats the one I'm struggling with the most??

OpenStudy (bradely):

1. f(x)=x+7 let y = x+7 y = x+7 x = y-7 f^(-1) (x) = x-7 = g(x) Both are inverse functions Step by step solution http://www.mathskey.com/question2answer/